OpenStudy (swaqqout_kid1):

Laura retired from her job recently, and she has saved about $377,957.00 over the course of her career. She plans to withdraw $1,358.00 each month to pay for living expenses. After a certain amount of time, the balance in Laura's account is $345,365.00. How many months have passed since Laura retired? @sammixboo

sammixboo (sammixboo):

Okie dokie so the equation we have is this \(\large \rm \color{#20bd23}{377,957~-~1,358m~=~345,365}\) We must first, add \(\rm \color{#20bd23}{377,957}\) on both sides of the equation... \(\large \rm \color{#20bd23}{1,358m~=~\color{red}{345,365~+~377,957}}\) Can you tell me what is \(\rm \color{red}{345,365~+~377,957}\)

OpenStudy (swaqqout_kid1):

723322

sammixboo (sammixboo):

Yes, it is \(\rm \color{#20bd23}{723,322}\), so now plug in \(\rm \color{#20bd23}{723,322}\) where \(\rm \color{red}{345,365~+~377,957}\) was in the equation \(\large \rm \color{#20bd23}{1,358m~=~723,322}\) Now we divide \(\rm \color{#20bd23}{1,358}\) on both sides of the equation \(\large \rm \color{red}{\frac{1,358m}{1,358}~=~\frac{723,322}{1,358}}\) \(\large \rm \color{#20bd23}{m~=~\color{red}{\frac{723,322}{1,358}}}\)

sammixboo (sammixboo):

So what is \(\rm \color{#20bd23}{723,322\div 1,358}\)

OpenStudy (swaqqout_kid1):

532

sammixboo (sammixboo):

Hold on a second let me check something

sammixboo (sammixboo):

Blah I am confusing myself maybe @sangya21 can help <_>

sammixboo (sammixboo):

I know the answer is 23-24 months, but I totally did the problem wrong, and I am not sure what I did. I am fixing to fall asleep as well

OpenStudy (anonymous):

\[$377,957.00 = ($1,358.00*m) + $345,365.00\] m = no. of months \[$377,957.00 - $345,365.00 = ($1,358.00*m) \] \[($377,957.00 - $345,365.00) /($1,358.00) = m \]

sammixboo (sammixboo):

OOOH I see what I did :P Anyways, night! I am going to bed

sammixboo (sammixboo):

I did the starting equation wrong